OpenStudy (anonymous):

perform the division. k^3+t^3/k+t

OpenStudy (anonymous):

\[\frac{ k ^{3}+t ^{3} }{ k+t }\]

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

\[\huge\rm (x^3 + y^3) = (x+y)(x^2 - xy + y^2)\] this is formula of cube factoring let me know if you need explanation of how its work

Nnesha (nnesha):

difference of cubes*

OpenStudy (anonymous):

 @Nnesha